OpenStudy (anonymous):

Write (6a)^2 without exponents (6a)^2 =

OpenStudy (anonymous):

6a(6a) , Iam not completely sure.

OpenStudy (anonymous):

I know theres some type of way or simple formula to solve this problem. I just forgot it and I can't find it in my notes. I'm looking for someone to tell me.

OpenStudy (kc_kennylau):

First you have to know that \(x^2=x\times x\). Therefore, \((6a)^2=6a\times6a\). Then you'll have to know that \(6a=6\times a\). Therefore, \((6a)^2=6a\times6a=6\times a\times6\times a=36a\times a\)